These equation has more than 2 solutions.
27x^3-8=0 Help. What do I do since it is a cube not square

Answer by midwood_trail(310)

27x^3 -8 = 0
Add 8 to both sides.
27x^3 = 8
Divide both sides by 27.
x^3 = (8/27)
We now take the CUBE ROOT of both sides.
Final answer is x = 2/3